Calculate the wavelength, in nanometers, of the spectra line produced when an electron in a hydrogen atom undergoes the transition from the energy level n 4 to the level n 2. Number nm
Answer #1

1/ λ = R * (1/(nf)^2 - 1/(ni)^2)

where R = Rydberg constant = 1.09737316 *10^7 m-1

nf = 2 and ni = 4

1/ λ = (1.09737316 *10^7 m^-1)*(1/2^2 - 1/4^2) = 2057574.68 m^-1

So the wave length is λ = 1/(2057574.68 m^-1) = 486 nm